# namerefThe definition of as changed

### v1

@@ VI Salient Features and Outlook
The main strength of our result is that the derived key rate under optimal collective attacks depends solely on two directly measurable quantities: the QBER and the CJWR steering violation $\mathcal{F}_{3}$. Similar to DI-QKD approaches based on Bell inequality violations [16], this avoids the need to reconstruct or assume an explicit quantum state model, allowing the key rate to be calculated from observed statistics. By reducing the analysis to an effective two-qubit Bell-diagonal scenario and relating $\mathcal{F}_{3}$ to Eve’s Holevo information, we obtain a closed-form expression for the asymptotic key rate that is both analytical and operationally meaningful. Whereas previous 1sDI-QKD protocols, like those by Branciard et al. [55], Tomamichel et al. [56], and Masini and Sarkar [58], rely upon entropy-based uncertainty relations or numerical postprocessing, our approach explicitly incorporates steering violation into the key rate bound. This makes the protocol particularly attractive for experimental implementations with constrained resources.
@@
Furthermore, our protocol remains robust against depolarizing noise, tolerating up to $8.62\%$ QBER by relying solely on the observed correlations $(Q,\mathcal{F}_{3})$. For comparison, DI-QKD protocols using three-setting Bell inequalities tolerate up to $Q_{c}=7.5\%$ [18], while those based on asymmetric Bell inequalities reach $Q_{c}=8.34\%$ [22]. The resulting performance approaches that of fully trusted DD schemes, yet preserves significant device-independence, reinforcing the value of 1sDI-QKD in practical implementations. Our protocol also shows favourable thresholds under detection inefficiency, remaining secure with efficiencies as low as $74.5\%$ under post-selection. This compares favorably to DI-QKD thresholds, which often demand efficiencies greater than $87\%$, as demonstrated in recent photonic DI-QKD (with noisy preprocessing) implementation [25].
@@
This work’s analytical framework opens up a number of promising directions for the advancement of 1sDI-QKD in practice. Incorporating noisy preprocessing [31, 73] is a crucial extension that could lower the critical detection-efficiency threshold and increase robustness against experimental imperfections. In addition, experimental studies  [59] show that linear steering inequalities can be modified to account for observed losses, as discussed in Section V. This suggests that our detection thresholds could be made more tolerant in practical implementations. Another significant step towards composable security is the establishment of finite-size security through the use of entropy accumulation techniques [67, 66]. Lastly, investigating alternative steering inequalities with improved loss and noise resilience [43, 44] may increase the applicability of our method across various quantum network architectures. Such directions collectively aim to strengthen the viability of 1sDI-QKD as a secure and scalable solution for near-term quantum communication.

### v2

@@ VI Salient Features and Outlook
A key strength of this work is the explicit analytical characterization of the asymptotic secret key rate under optimal collective attacks. By relating the experimentally observed CJWR steering violation $\mathcal{F}_{3}$ to Eve’s Holevo information within an effective two-qubit Bell-diagonal description, we derive a closed-form Devetak–Winter key-rate expression that can be evaluated directly from measured statistics. In contrast to QBER-based approaches, our protocol follows an E91-style logic, analogous to device-independent QKD based on Bell-inequality violations [16], in which Eve’s information is quantified directly through the degree of CJWR steering violation.
@@
Furthermore, our protocol remains robust against depolarizing noise, tolerating up to $8.62\%$ QBER by relying solely on the observed correlations $(Q,\mathcal{F}_{3})$. For comparison, DI-QKD protocols using three-setting Bell inequalities tolerate up to $Q_{c}=7.5\%$ [18], while those based on asymmetric Bell inequalities reach $Q_{c}=8.34\%$ [22]. The resulting performance approaches that of fully trusted DD schemes, yet preserves significant device-independence, reinforcing the value of 1sDI-QKD in practical implementations. Our protocol also shows favourable thresholds under detection inefficiency, remaining secure with efficiencies as low as $74.5\%$ under post-selection. This compares favourably to DI-QKD thresholds, which often demand efficiencies greater than $87\%$, as demonstrated in recent photonic DI-QKD (with noisy preprocessing) implementation [25].
While the present analysis is restricted to asymptotic security against collective attacks, such attacks capture the optimal single-round eavesdropping strategy and have historically played a central role in the development of device-independent security proofs [16]. In standard, device-dependent scenarios, collective-attack bounds can often be lifted to coherent-attack security via de Finetti arguments; however, in DI and 1sDI settings, this route is considerably more involved, as known quantum de Finetti theorems rely on assumptions such as bounded local dimension that are not available when devices are uncharacterized [70]. A full extension to composable security against general coherent attacks, including finite-size effects, therefore requires alternative tools, most notably entropy accumulation techniques capable of handling general sequential and memory-dependent strategies [63, 64]. In this context, we expect that the explicit single-round bounds on Eve’s information derived here will provide a natural and valuable building block for such future extensions beyond the collective-attack regime.
@@
This work’s analytical framework opens up a number of promising directions for the advancement of 1sDI-QKD in practice. Incorporating noisy preprocessing [31, 71] is a crucial extension that could lower the critical detection-efficiency threshold and increase robustness against experimental imperfections. In addition, experimental studies  [56] show that linear steering inequalities can be modified to account for observed losses, as discussed in Section V. This suggests that our detection thresholds could be made more tolerant in practical implementations. Lastly, investigating alternative steering inequalities with improved loss and noise resilience [43, 44] may increase the applicability of our method across various quantum network architectures. Such directions collectively aim to strengthen the viability of 1sDI-QKD as a secure and scalable solution for near-term quantum communication.